# Out-of-Distribution Detection by Leveraging Between-Layer Transformation Smoothness

### v1

@@ Abstract
Effective OOD detection is crucial for reliable machine learning models, yet most current methods are limited in practical use due to requirements like access to training data or intervention in training. We present a novel method for detecting OOD data in deep neural networks based on transformation smoothness between intermediate layers of a network (BLOOD), which is applicable to pre-trained models without access to training data. BLOOD utilizes the tendency of between-layer representation transformations of in-distribution (ID) data to be smoother than the corresponding transformations of OOD data, a property that we also demonstrate empirically for Transformer networks. We evaluate BLOOD on several text classification tasks with Transformer networks and demonstrate that it outperforms methods with comparable resource requirements. Our analysis also suggests that when learning simpler tasks, OOD data transformations maintain their original sharpness, whereas sharpness increases with more complex tasks.
@@ 1 Introduction
Owing to their consistent state-of-the-art performance across diverse ML tasks (Abiodun et al., 2018), Deep Neural Networks (DNNs) have garnered significant attention in OOD detection research. While popular baselines make use of the model’s posterior class probabilities (Hendrycks & Gimpel, 2017), the issue of overconfidence in DNNs (Guo et al., 2017) frequently erodes the credibility of these probabilities. An alternative is offered by the group of methods that leverage the fundamental concept in deep neural networks, namely, representation learning. Because a DNN encodes similar instances closely in its representation space, an OOD instance can be identified based on the distance between its representation and the representations of other instances in the training set (Lee et al., 2018). The downside of these methods, however, is that they require the presence of training data during prediction or involve intervention in the model’s training procedure. This is a significant practical limitation, as using third-party models pre-trained on non-public data is increasingly the standard practice. A case in point is the Hugging Face Transformers library (Wolf et al., 2020), which provides readily available community models but often lacks comprehensive training data or detailed training procedures.
@@
In this paper, we introduce a novel OOD detection method that leverages the inherent differences in how DNNs process ID and OOD data. The method is white-box and has the potential for broad practical applicability. More concretely, our Between Layer Out-Of-Distribution (BLOOD) Detection method estimates the smoothness of between-layer transformations of intermediate representation, building on the insight that these transformations tend to be smoother for ID data than for OOD data. We evaluate BLOOD on Transformer-based (Vaswani et al., 2017) pre-trained large language models applied to text classification, the most prevalent task in natural language processing (NLP), and find that it outperforms other state-of-the-art OOD detection white-box methods and even some open-box methods. We further analyze BLOOD to probe into the underlying causes of the differences between how ID and OOD intermediate representations are transformed and evaluate BLOOD on two other types of distribution shifts – semantic and background shift. We provide code and data used in our experiments for result reproducibility. [^2] https://github.com/fjelenic/between-layer-ood
@@ 2 Related Work
Black-box.  Methods with minimal prerequisites typically rely on posterior class probabilities, operating under the assumption that when a model exhibits uncertainty about an instance, the instance is more likely to be OOD. A commonly used baseline quantifies the uncertainty of an instance as the negative of the model’s maximum softmax probability for that instance (Lee et al., 2018). A straightforward modification employs the entropy of softmax probabilities rather than the maximum value. Liu et al. (2020b) proposed using energy scores instead of softmax scores to overcome the issue of DNN overconfidence.
@@
White-box.  Gal & Ghahramani (2016) proposed using Monte-Carlo dropout to more reliably estimate the model’s uncertainty, showing that dropout (Srivastava et al., 2014) with DNNs approximates Bayesian inference. Although Monte-Carlo dropout outperforms vanilla posterior probabilities in OOD detection (Ovadia et al., 2019), it is computationally expensive as it requires multiple forward passes. Another way of leveraging the access to model’s architecture is to use gradients to implicitly measure the uncertainty of the model’s predictions (Oberdiek et al., 2018; Huang et al., 2021). Gradient methods primarily employ the gradient norm to gauge the difference between the model’s posterior distribution and the ideal distribution.
@@
Open-box.  Because DNNs posterior probabilities tend to exhibit overconfidence, Guo et al. (2017) suggested using temperature scaling to calibrate the model’s posterior probabilities, which entails the usage of a separate validation set. To get higher quality predictive uncertainty estimates, Lakshminarayanan et al. (2017) train an ensemble of differently initialized models and combine their predictions. Although ensembles are robust to different distributional shifts (Ovadia et al., 2019), they impose a significant computational and memory overhead because they require training and keeping in memory of multiple models. A popular approach to OOD detection for DNNs revolves around the utilization of information related to distances in the representation space (Lee et al., 2018; Van Amersfoort et al., 2020; Liu et al., 2020a; Hsu et al., 2020; Kuan & Mueller, 2022; Sun et al., 2022). However, these approaches require access to the training data or changes in the standard training procedure. Yet another set of methods relies on exposing the model to OOD samples during training to improve the performance on OOD detection task (Hendrycks et al., 2019; Thulasidasan et al., 2021; Roy et al., 2022). Still, a major practical limitation of these methods is the necessity for OOD data, whose entire distribution is typically unknown in real-world scenarios. Several post-hoc methods also need OOD data, but for validation sets to optimize their method’s hyperparameters (Liang et al., 2018; Sun et al., 2021; Sun & Li, 2022).
@@ 3.2 Intuition
DNNs work by mapping the input features onto a high-dimensional representation space through $L$ layers, creating a representation of the data suitable for the task at hand. The mapping is realized as a composition of several non-linear functions, where each function creates an intermediate representation of the input. Representations of DNNs with useful inductive biases for the task at hand tend to gradually progress from input features towards more abstract representation levels through layers, i.e., lower layers model lower-level features while upper layers model higher-level features. For example, when convolutional neural networks (CNNs) process images, they can first represent the simplest features in an image, e.g., lines and edges, which are then used to create representations of textures and simple shapes that are then combined to represent objects (LeCun et al., 1998). Recently, Vision Transformers (ViT) (Dosovitskiy et al., 2021), which are garnering popularity in computer vision, were shown to process data in a similar fashion (Ghiasi et al., 2022). Likewise, Peters et al. (2018); Tenney et al. (2019); Jawahar et al. (2019) showed that large transformer-based language models create text representations that progress gradually from representations that encode morphological and syntactic information at the lower DNN layers to representations that encode semantic meaning in the upper layers.
We hypothesize that during the model’s training, the model learns smooth transformations between DNN layers corresponding to natural and meaningful progressions between abstractions for ID data. We further hypothesize that these progressions will not match OOD data, hence the transformations will not be smooth for OOD data. Thus, if we could measure the smoothness of transformations in representations between layers, we could in principle differentiate between ID and OOD data. We also speculate that the difference in smoothness of transformations between ID and OOD data should be emphasized in the upper layers of DNNs. Lower layers typically represent low-level features that are more universal, whereas upper layers tend to cluster instances around task-specific features that are not shared between ID and OOD data, potentially creating a mismatch in levels of abstraction.
@@ 4.1 Experimental setup
We compare BLOOD to several state-of-the-art black-box and white-box OOD detection methods: (1) Maximum softmax probability(MSP) – the negative posterior class probability of the most probable class, $-\max_{c}p(y=c|{\bm{x}})$, often considered a baseline OOD detection method(Hendrycks & Gimpel, 2017); (2) Entropy(ENT) – the entropy of the posterior class distribution, ${\mathbb{H}}[Y|{\bm{x}},{\bm{w}}]$; (3) Energy(EGY) – a density-based method that overcomes the overconfidence issue by calculating energy scores from logits $-\log\sum_{i=0}^{C-1}e^{f_{L}({\bm{x}})_{i}}$ instead of softmax scores (Liu et al., 2020b); (4) Monte-Carlo dropout(MC) – the entropy of predictive distribution obtained using Monte-Carlo dropout (Gal & Ghahramani, 2016). We use $M=30$ stochastic forward passes to estimate uncertainty; (5) Gradient norm(GRAD) – the L2-norm of the penultimate layer’s gradient of the loss function with most likely class considered as a true class (Oberdiek et al., 2018).
Additionally, we compare BLOOD to three standard open-box OOD detection methods. Given that these methods entail considerably more prerequisites compared to BLOOD and other white/black-box methods, this comparison is intended solely as a reference point: (1) Ensemble(ENSM) – an ensemble of $M=5$ models of the same type, e.g., an ensemble of five RoBERTa or ensemble of five ELECTRA models, (Lakshminarayanan et al., 2017); (2) Temperature scaling(TEMP) – introduces a temperature parameter $T$ into the softmax function such that it minimizes the negative log-likelihood on the ID validation set (Guo et al., 2017); (3) Mahalanobis distance(MD) – Mahalanobis distance of a query instance in the representation space with respect to the closest class-conditional Gaussian distribution (Lee et al., 2018).
@@ 4.2 OOD detection performance
As the performance measure for OOD detection, we follow the standard practice and use the area under the receiver operating characteristic curve (AUROC) metric (in Appendix F, we report the results using two other commonly used metrics, AUPR-IN and FPR@95TPR; these gave qualitatively identical results as AUROC). The OOD detection task is essentially a binary classification task, with AUC corresponding to the probability that a randomly chosen OOD instance will have a higher uncertainty score than a randomly chosen ID instance (Fawcett, 2006). The AUROC for random value assignment is 50%, while a perfect method achieves 100%. We run each experiment five times with different random seeds and report the mean AUROC.
@@
RoBERTa sst 50.56 72.83 71.69 71.69 71.61 68.28 71.76 69.03 71.64 85.36
subj 52.02 74.66 74.55 74.55 75.79 74.21 74.93 76.68 74.41 93.47
agn 77.46 61.95 73.57 73.80 76.36 77.55 73.58 80.35 75.38 82.63
trec 69.63 95.30 96.20 96.40 96.28 95.68 96.14 96.87 96.74 96.74
bp $*{}^{*}$ $*{}^{*}$ 70.15 72.82 85.84 74.29 73.11 79.39 86.01 97.35
ar $*{}^{*}$ $*{}^{*}$ 89.06 89.96 92.39 90.59 88.65 92.44 92.25 98.35
mg $*{}^{*}$ $*{}^{*}$ 75.02 76.60 86.45 79.98 74.28 76.98 84.30 95.12
ng $*{}^{*}$ 72.02 77.49 78.76 82.65 79.32 76.93 80.77 82.87 90.68
ELECTRA sst 74.36 $*{}^{*}$ 73.84 73.84 71.97 70.81 73.82 73.81 73.58 78.85
subj 74.10 77.41 78.17 78.17 70.46 77.71 78.11 79.23 78.20 81.59
agn 65.67 80.28 76.80 77.01 79.75 79.55 76.57 79.50 78.31 86.10
trec 97.48 $*{}^{*}$ 97.26 97.56 97.48 96.21 97.07 97.55 98.20 97.54
bp $*{}^{*}$ $*{}^{*}$ 78.56 81.75 84.63 83.04 76.77 84.20 84.69 98.28
ar 84.58 $*{}^{*}$ 87.74 88.44 90.64 88.53 87.52 91.98 90.35 95.47
mg 80.52 $*{}^{*}$ 73.83 74.78 80.41 76.67 73.35 76.86 78.47 92.96
ng 77.61 $*{}^{*}$ 76.45 77.73 80.83 79.11 75.97 79.93 80.75 89.13
@@ 4.3 Source of the differences in transformations of ID and OOD data
Table 2 shows the effect size quantified using CLES for the changes in representations between ID and OOD data. In most setups, CLES is far above 50%, which means that representations of ID data underwent more significant changes than those of OOD data. The results imply that the learning algorithm’s focus during training is on the ID region of the representation space. In contrast, the rest of the representation space is largely unaltered. Furthermore, the difference in transformation smoothness between layers, observed between ID and OOD data, can be attributed to the inherently non-smooth transformations of the initalized models. These non-smooth transformations gradually become smoother within the ID region. However, more complex datasets (bp, ar, mg, and ng) in conjunction with the RoBERTa model contradict our initial expectation. In these cases, CLES approaches or even drops below 50%. This indicates that the ID region of the representation space undergoes similar or even fewer changes compared to the rest of the representation space.
@@ 4.4 The effect of dataset complexity
Table 3: The performance of OOD detection methods for the simplified datasets measured by AUROC (%). The best-performing white/black-box method is in bold. Open-box methods that outperform all white/black-box methods are in bold. Higher is better. Right side of the table shows a comparison of changes in representations between ID and OOD data using CLES (%).
@@
We suspect that the increase in AUROC for the other white/black-box methods may be attributed to the same factor that led to the AUROC decrease in BLOOD – namely, the task’s simplicity. However, this cause manifests differently. The simplified datasets, having fewer ambiguous instances in their test sets due to the reduced number of classes, allow the other (probabilistic) methods to more accurately attribute the estimated uncertainty to the OOD data. See Appendix D for a more detailed explanation and visualization using dataset cartography (Swayamdipta et al., 2020).

### v2

@@ Abstract
Effective out-of-distribution (OOD) detection is crucial for reliable machine learning models, yet most current methods are limited in practical use due to requirements like access to training data or intervention in training. We present a novel method for detecting OOD data in Transformers based on transformation smoothness between intermediate layers of a network (BLOOD), which is applicable to pre-trained models without access to training data. BLOOD utilizes the tendency of between-layer representation transformations of in-distribution (ID) data to be smoother than the corresponding transformations of OOD data, a property that we also demonstrate empirically. We evaluate BLOOD on several text classification tasks with Transformer networks and demonstrate that it outperforms methods with comparable resource requirements. Our analysis also suggests that when learning simpler tasks, OOD data transformations maintain their original sharpness, whereas sharpness increases with more complex tasks.
@@ 1 Introduction
Owing to their consistent state-of-the-art performance across diverse ML tasks (Abiodun et al., 2018), Deep Neural Networks (DNNs) have garnered significant attention in OOD detection research. While popular baselines make use of the model’s posterior class probabilities (Hendrycks & Gimpel, 2017), the issue of overconfidence in DNNs (Guo et al., 2017) frequently erodes the credibility of these probabilities. An alternative is offered by the group of methods that leverage the fundamental concept of DNNs, namely, representation learning. Because a DNN encodes similar instances closely in its representation space, an OOD instance can be identified based on the distance between its representation and the representations of other instances in the training set (Lee et al., 2018). The downside of these methods, however, is that they require the presence of training data during prediction or involve intervention in the model’s training procedure. This is a significant practical limitation, as using third-party models pre-trained on non-public data is increasingly the standard practice. A case in point is the Hugging Face Transformers library (Wolf et al., 2020), which provides community models but often lacks comprehensive details about their training.
@@
In this paper, we focus on the OOD detection for the Transformer architecture (Vaswani et al., 2017), which has emerged as the predominant architecture in numerous ML domains. We introduce a novel OOD detection method that leverages the inherent differences in how Transformers process ID and OOD data. The method is white-box and has the potential for broad practical applicability. More concretely, our Between Layer Out-Of-Distribution (BLOOD) Detection method estimates the smoothness of between-layer transformations of intermediate representation, building on the insight that these transformations tend to be smoother for ID data than for OOD data. We evaluate BLOOD on Transformer-based pre-trained large language models applied to text classification, the most prevalent task in natural language processing (NLP), and find that it outperforms other state-of-the-art OOD detection white-box methods and even some open-box methods. We further analyze BLOOD to probe into the underlying causes of the differences between how ID and OOD intermediate representations are transformed and evaluate BLOOD on two other types of distribution shifts – semantic and background shift. We provide code and data for our experiments. [^2] https://github.com/fjelenic/between-layer-ood
@@ 2 Related Work
Black-box.  Methods with minimal prerequisites typically rely on posterior class probabilities, assuming that when a model is uncertain about an instance, the instance is more likely to be OOD. A commonly used baseline quantifies the uncertainty of an instance as the negative of the model’s maximum softmax probability for that instance (Lee et al., 2018). A straightforward modification employs the entropy of softmax probabilities rather than the maximum value. Liu et al. (2020b) proposed using energy scores instead of softmax scores to overcome the issue of DNN overconfidence.
@@
White-box.  Gal & Ghahramani (2016) proposed using Monte-Carlo dropout to more reliably estimate the model’s uncertainty, showing that dropout (Srivastava et al., 2014) with DNNs approximates Bayesian inference. Although Monte-Carlo dropout outperforms vanilla posterior probabilities in OOD detection (Ovadia et al., 2019), it is computationally expensive as it requires multiple forward passes. Another way of leveraging the access to model’s architecture is to use gradients to implicitly measure the uncertainty of the model’s predictions (Oberdiek et al., 2018; Huang et al., 2021). Gradient methods primarily employ the gradient norm to gauge the difference between the model’s posterior distribution and the ideal distribution. Djurisic et al. (2023) detect OOD data by pruning and adjusting the representations of the model, grounded in the intuition that the representations generated by contemporary DNNs tend to be excessive for their designated tasks.
@@
Open-box.  Because DNNs posterior probabilities tend to exhibit overconfidence, Guo et al. (2017) suggested using temperature scaling to calibrate the model’s posterior probabilities, which entails the usage of a separate validation set. To get higher quality predictive uncertainty estimates, Lakshminarayanan et al. (2017) train an ensemble of differently initialized models and combine their predictions. Although ensembles are robust to different distributional shifts (Ovadia et al., 2019), they impose a significant computational and memory overhead because they require training and keeping in memory of multiple models. Agarwal et al. (2022) extend the gradient-based methods by leveraging the variance of the gradient of the predicted label w.r.t. the input through different checkpoints during training. A popular approach to OOD detection for DNNs revolves around the utilization of information related to distances in the representation space (Lee et al., 2018; Van Amersfoort et al., 2020; Liu et al., 2020a; Hsu et al., 2020; Kuan & Mueller, 2022; Sun et al., 2022). However, these approaches require access to the training data or changes in the standard training procedure. Yet another set of methods relies on exposing the model to OOD samples during training to improve the performance on OOD detection task (Hendrycks et al., 2019; Thulasidasan et al., 2021; Roy et al., 2022). Still, a major practical limitation of these methods is the necessity for OOD data, whose entire distribution is typically unknown in real-world scenarios. Several post-hoc methods also need OOD data, but for validation sets to optimize their method’s hyperparameters (Liang et al., 2018; Sun et al., 2021; Sun & Li, 2022).
@@ 3.2 Intuition
Transformers work by mapping the input features onto a high-dimensional representation space through $L$ layers using the self-attention mechanism, creating a representation of the data suitable for the task at hand. The mapping is realized as a composition of several attention layers, where each layer creates an intermediate representation of the input. It has been show that Transformer-based models tend to gradually progress from input features towards more abstract representation levels through layers, i.e., lower layers model lower-level features, while upper layers model higher-level features. For example, Peters et al. (2018); Tenney et al. (2019); Jawahar et al. (2019) showed that large Transformer-based language models create text representations that progress gradually from representations that encode morphological and syntactic information at the lower layers to representations that encode semantic meaning in the upper layers. Likewise, Vision Transformers (ViT) (Dosovitskiy et al., 2021), which are garnering popularity in computer vision, were shown to process images in a similar fashion (Ghiasi et al., 2022).
We hypothesize that during the model’s training, the model learns smooth transformations between layers corresponding to natural and meaningful progressions between abstractions for ID data. We further hypothesize that these progressions will not match OOD data, hence the transformations will not be smooth for OOD data. Thus, if we could measure the smoothness of transformations in representations between layers, we could in principle differentiate between ID and OOD data. We also speculate that the difference in smoothness of transformations between ID and OOD data should be emphasized in the upper layers of a Transformer. Lower layers typically represent low-level features that are more universal, whereas upper layers tend to cluster instances around task-specific features that are not shared between ID and OOD data, potentially creating a mismatch in levels of abstraction.
@@ 4.1 Experimental setup
We compare BLOOD to several state-of-the-art black-box and white-box OOD detection methods: (1) Maximum softmax probability(MSP) – the negative posterior class probability of the most probable class, $-\max_{c}p(y=c|{\bm{x}})$, often considered a baseline OOD detection method(Hendrycks & Gimpel, 2017); (2) Entropy(ENT) – the entropy of the posterior class distribution, ${\mathbb{H}}[Y|{\bm{x}},{\bm{w}}]$; (3) Energy(EGY) – a density-based method that overcomes the overconfidence issue by calculating energy scores from logits $-\log\sum_{i=0}^{C-1}e^{f_{L}({\bm{x}})_{i}}$ instead of softmax scores (Liu et al., 2020b); (4) Monte-Carlo dropout(MC) – the entropy of predictive distribution obtained using Monte-Carlo dropout (Gal & Ghahramani, 2016). We use $M=30$ stochastic forward passes to estimate uncertainty; (5) Gradient norm(GRAD) – the L2-norm of the penultimate layer’s gradient of the loss function with most likely class considered as a true class (Oberdiek et al., 2018). (6) Activation shaping(ASH) – removing 90% of the smallest activations and adjusting the rest using ASH-S method in the penultimate layer (Djurisic et al., 2023).
Additionally, we compare BLOOD to three standard open-box OOD detection methods. Given that these methods entail considerably more prerequisites compared to BLOOD and other white/black-box methods, this comparison is intended solely as a reference point: (1) Rectified Activations(ReAct) – setting the values of the activations in the penultimate layer to be at most the 90th percentile of the activations of the training data (Sun et al., 2021). (2) Ensemble(ENSM) – an ensemble of $M=5$ models of the same type, e.g., an ensemble of five RoBERTa or ensemble of five ELECTRA models, (Lakshminarayanan et al., 2017); (3) Temperature scaling(TEMP) – introduces a temperature parameter $T$ into the softmax function such that it minimizes the negative log-likelihood on the ID validation set (Guo et al., 2017); (4) Mahalanobis distance(MD) – Mahalanobis distance of a query instance in the representation space with respect to the closest class-conditional Gaussian distribution (Lee et al., 2018).
@@ 4.2 OOD detection performance
As the performance measure for OOD detection, we follow the standard practice and use the area under the receiver operating characteristic curve (AUROC) metric (in Appendix H, we report the results using two other commonly used metrics, AUPR-IN and FPR@95TPR; these gave qualitatively identical results as AUROC). The OOD detection task is essentially a binary classification task, with AUC corresponding to the probability that a randomly chosen OOD instance will have a higher uncertainty score than a randomly chosen ID instance (Fawcett, 2006). The AUROC for random value assignment is 50%, while a perfect method achieves 100%. We run each experiment five times with different random seeds and report the mean AUROC.
@@
RoBERTa sst 50.56 72.83 71.69 71.69 71.61 68.28 71.76 67.22 69.55 69.03 71.64 85.36
subj 52.02 74.66 74.55 74.55 75.79 74.21 74.93 79.27 73.33 76.68 74.41 93.47
agn 77.46 61.95 73.57 73.80 76.36 77.55 73.58 72.54 77.10 80.35 75.38 82.63
trec 69.63 95.30 96.20 96.40 96.28 95.68 96.14 90.36 96.05 96.87 96.74 96.74
bp $*{}^{*}$ $*{}^{*}$ 70.15 72.82 85.84 74.29 73.11 82.18 86.19 79.39 86.01 97.35
ar $*{}^{*}$ $*{}^{*}$ 89.06 89.96 92.39 90.59 88.65 91.42 92.65 92.44 92.25 98.35
mg $*{}^{*}$ $*{}^{*}$ 75.02 76.60 86.45 79.98 74.28 81.62 87.30 76.98 84.30 95.12
ng $*{}^{*}$ 72.02 77.49 78.76 82.65 79.32 76.93 77.73 83.17 80.77 82.87 90.68
ELECTRA sst 74.36 $*{}^{*}$ 73.84 73.84 71.97 70.81 73.82 67.92 71.18 73.81 73.58 78.85
subj 74.10 77.41 78.17 78.17 70.46 77.71 78.11 75.11 68.33 79.23 78.20 81.59
agn 65.67 80.28 76.80 77.01 79.75 79.55 76.57 77.96 79.46 79.50 78.31 86.10
trec 97.48 $*{}^{*}$ 97.26 97.56 97.48 96.21 97.07 90.18 97.50 97.55 98.20 97.54
bp $*{}^{*}$ $*{}^{*}$ 78.56 81.75 84.63 83.04 76.77 79.81 85.26 84.20 84.69 98.28
ar 84.58 $*{}^{*}$ 87.74 88.44 90.64 88.53 87.52 83.96 91.01 91.98 90.35 95.47
mg 80.52 $*{}^{*}$ 73.83 74.78 80.41 76.67 73.35 71.84 81.22 76.86 78.47 92.96
ng 77.61 $*{}^{*}$ 76.45 77.73 80.83 79.11 75.97 74.50 80.95 79.93 80.75 89.13
@@ 4.3 Source of the differences in transformations of ID and OOD data
Table 2 shows the effect size quantified using CLES for the changes in representations between ID and OOD data. In most setups, CLES is far above 50%, which means that representations of ID data underwent more significant changes than those of OOD data. The results imply that the learning algorithm’s focus during training is on the ID region of the representation space. In contrast, the rest of the representation space is largely unaltered. Moreover, the difference in transformation smoothness between layers, observed between ID and OOD data, can be attributed to the inherently non-smooth transformations of the initialized models. These non-smooth transformations gradually become smoother within the ID region. However, more complex datasets (bp, ar, mg, and ng) in conjunction with the RoBERTa model contradict our initial expectation. In these cases, CLES approaches or even drops below 50%. This indicates that the ID region of the representation space undergoes similar or even lesser changes compared to the rest of the representation space.
@@ 4.4 The effect of dataset complexity
Table 3: The performance of OOD detection methods for the simplified datasets measured by AUROC (%). The best-performing white/black-box method is in bold. Open-box methods that outperform all white/black-box methods are in bold. Higher is better. The right side of the table shows a comparison of changes in representations between ID and OOD data using CLES (%).
@@
We suspect that the increase in AUROC for the other white/black-box methods may be attributed to the same factor that led to the AUROC decrease in BLOOD – namely, the task’s simplicity. However, this cause manifests differently. The simplified datasets, having fewer ambiguous instances in their test sets due to the reduced number of classes, allow the other (probabilistic) methods to more accurately attribute the estimated uncertainty to the OOD data. See Appendix F for a more detailed explanation and visualization using dataset cartography (Swayamdipta et al., 2020).